# Pruning the Unsurprising: Efficient Code Reasoning via First-Token Surprisal

### v1

@@ 3.2 Anchor-guided Pruning
Specifically, we design a pattern-matching algorithm that verifies whether each step in $C_{coarse}$ corresponds to a matching step in $C$ while preserving their original order. The matching is performed using Gestalt Pattern Matching (Black 2004) as a text similarity metric. A pruning is considered valid only if all steps in $C_{coarse}$ achieve a similarity score above a predefined threshold $\tau$ when matched against sequential steps in $C$. This guarantees that $C_{coarse}$ forms a consistent substructure of the original reasoning trace. The full pattern-matching algorithm is detailed in Algorithm 1. We leverage high-temperature sampling, which provides the necessary diversity to efficiently re-prompt failed cases, ensuring that a valid $C_{\text{coarse}}$ can be eventually generated.
@@
3.3 Surprisal-based Refining
Following the coarse-grained pruning, we perform a meticulous, logic-aware refinement in $C_{coarse}$ to identify more subtle redundancies within the core reasoning path.
First-Token Surprisal as Logical Importance
We introduce First-Token Surprisal as a novel metric to precisely quantify the logical importance of each step, enabling us to iteratively filter out the least informative ones and produce the final highly condensed CoT.
In information theory, entropy is used to quantify the average level of uncertainty or information associated with a variable’s potential states or possible outcomes (Shannon 1948). In the era of LLMs, this concept can be applied to each token generated by the model. Given an LLM generates a sequence of tokens $x_{1},x_{2},\dots,x_{N}$, the conditional probability of the $n$-th token is given by $p(x_{n}|x_{<n})$, derived from the model’s logits. Token-level entropy measures the model’s uncertainty over the entire vocabulary $V$ for the next token (Malinin and Gales 2020; Kuhn, Gal, and Farquhar 2023). It is the surprisal over all possible outcomes:
$H(x_{n}|x_{<n})=-\sum_{v\in V}p(v|x_{<n})\log p(v|x_{<n})$ (1)
In large reasoning models, the first token of each step often signals a logical transition or continuation. High entropy indicates the model’s uncertainty, suggesting that the step may introduce semantically or logically significant information. Conversely, low entropy typically reflects predictable continuations or syntactic fillers, implying limited contribution to the overall reasoning process (Wang et al. 2025; Cheng et al. 2025).
However, since the actual next token is known in the sequence, we directly compute the surprisal of the target token, a more efficient and targeted metric that quantifies the unexpectedness of the target token $x_{t}$:
$S(x_{t}|x_{<n})=-\log p(x_{t}|x_{<n})$ (2)
Surprisal quantifies how unexpected a given token is to the model. We focus specifically on the first-token surprisal of each reasoning step. A high surprisal value at the beginning of a step indicates logical salience—that is, the step either initiates a new line of reasoning or introduces novel information. In contrast, a low surprisal value suggests a continuation of prior context or redundant elaboration. Accordingly, we prune low-surprisal steps to obtain more concise and informative reasoning traces.
Iterative Pruning Algorithm
Based on the first-token surprisal metric, we propose a length-controllable iterative pruning algorithm. As shown in Algorithm 2, it calculates the surprisal score for the first token of each reasoning step within the $C_{coarse}$. It then iteratively removes the step with the lowest surprisal score until the total length of the CoT falls below a predefined token budget $L_{max}$. This process yields the final $C^{\prime}$, which is highly compressed while retaining the most logically significant steps.

### v2

@@ 3.2 Anchor-guided Pruning
Crucially, to mitigate LLM hallucination during compression, we enforce an extractive constraint, which validates structural and semantic alignment with $C$. Specifically, we design a pattern-matching algorithm that verifies whether each step in $C_{coarse}$ corresponds to a matching step in $C$ while preserving their original order. The matching is performed using Gestalt Pattern Matching Black (2004) as a text similarity metric. A pruning is considered valid only if all steps in $C_{coarse}$ achieve a similarity score above a predefined threshold $\tau$ when matched against sequential steps in $C$. The full pattern-matching algorithm is detailed in Algorithm 1 (see Appendix A). We leverage high-temperature sampling, which provides the necessary diversity to efficiently re-prompt failed cases, ensuring that a valid $C_{\text{coarse}}$ can be eventually generated.
3.3 Surprisal-based Refining
Following the coarse-grained pruning, the resulting $C_{coarse}$ may still contain verbose steps that contribute little to the logic. Grounded in our empirical finding that logical information is concentrated (Section 2), we perform a meticulous, logic-aware refinement in $C_{coarse}$ to identify more subtle redundancies within the core reasoning path.
First-Token Surprisal as Logical Importance.
We introduce First-Token Surprisal as a novel metric to precisely quantify the logical importance of each step, enabling us to filter out the least informative ones and produce the final highly condensed CoT. Let a reasoning step $s$ be a sequence of tokens $s=(x_{1},x_{2},\dots,x_{T})$. The informational value of $s$ within the context of previous steps $\mathcal{C}_{pre}$ is typically estimated by its joint probability. However, our analysis reveals that the first token $x_{1}$ serves as the “cognitive pivot” carrying the majority of the uncertainty. Therefore, we define the First-Token Surprisal $\mathcal{S}(s)$ as:
$\mathcal{S}(s\mid\mathcal{C}_{pre})=-\log P_{\theta}(x_{1}\mid\mathcal{C}_{pre})$ (1)
where $P_{\theta}$ denotes the probability distribution of the LRM. A high $\mathcal{S}(s)$ indicates a high-information transition (e.g., initiating a new deduction or self-correction), whereas a low score suggests a deterministic continuation or syntactic filler.
Pruning using First-Token Surprisal.
We formulate the fine-grained pruning as a constrained maximization problem. Our goal is to select a subset of steps $S^{\prime}\subset C_{coarse}$ that maximizes the total logical information subject to a length budget $L_{max}$:
$\begin{split}S^{*}=\operatorname*{arg\,max}_{S^{\prime}\subseteq C_{coarse}}\sum_{s\in S^{\prime}}\mathcal{S}(s)\\ \text{s.t.}\quad\sum_{s\in S^{\prime}}\text{len}(s)\leq L_{max}\end{split}$ (2)
This formulation explicitly prioritizes steps with high information density. To solve this efficiently, we employ a greedy iterative strategy. We calculate the surprisal score for all steps in $C_{coarse}$ and iteratively remove the step with the lowest $\mathcal{S}(s)$, while the relative order of steps in $S^{\prime}$ is preserved. The detailed procedure is provided in Algorithm 2 (see Appendix A). This process yields the final fine-grained CoT $C^{\prime}$, which retains the critical “aha moments” Guo et al. (2025) while meeting efficiency constraints.
3.4 Supervised Fine-tuning
Following the pruning, we construct the final training dataset $\mathcal{D}^{\prime}=\{(Q_{i},\mathbf{C}_{i}^{\prime},A_{i})\}_{i=1}^{N}$. For each instance, we concatenate the pruned CoT ($\mathbf{C}_{i}^{\prime}$) and the final answer ($A_{i}$) to form the complete target response $R_{i}$. We then fine-tune the LRM to minimize the standard negative log-likelihood of the target response tokens, conditioned on the input question. Formally, the loss is defined as:
where $r_{i,j}$ is the $j$-th token of the target response $R_{i}$, and $\theta$ represents the parameters of the model being fine-tuned. This supervised fine-tuning process effectively distills the knowledge from our pruning framework into the model. By training on these compact, logically salient examples, the model learns to internalize efficient reasoning patterns.
4 Experiments
Methods HE+ LCBv1_v3 LCBv4_v5 LCD
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 68.29 3051 1.16 42.16 7088 3.59 25.37 8336 5.15 19.74 8680 4.95
Original 75.61 2973 1.12 52.12 6611 3.15 30.97 8289 4.83 25.00 8485 4.72
Selective Context 54.88 2979 1.13 30.23 7025 3.75 16.79 8558 5.35 15.79 8461 4.90
LLMLingua-2 68.29 3075 1.19 38.89 6953 3.60 22.76 8474 5.31 17.54 8513 4.81
TokenSkip 73.78 2823 1.07 32.35 7095 3.85 20.15 8400 5.37 18.42 8503 4.87
SPIRIT 75.61 2764 1.07 50.82 6524 3.09 33.58 7892 4.62 25.00 8186 4.45
ASAP 78.66 2464 0.98 54.74 5177 2.09 36.19 6035 2.61 27.63 7541 3.48
Table 1: Experimental results of different methods on code generation benchmarks with DeepSeek-R1-Distill-Qwen-7B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds. The best results are highlighted in bold, and the second-best are underlined.
Methods GSM8K MATH500 AIME24 AIME25
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 83.55 1301 0.27 60.40 2629 0.70 36.67 8352 6.76 40.00 8145 6.67
Original 86.35 1250 0.26 63.80 2511 0.66 46.67 8034 6.43 43.33 8026 6.75
Selective Context 75.44 1108 0.24 52.20 2507 0.66 16.67 9610 7.40 10.00 9329 7.40
LLMLingua-2 79.98 1128 0.24 54.60 2802 0.76 36.67 8369 6.60 23.33 8919 7.42
TokenSkip 85.37 1303 0.27 65.60 2483 0.65 36.67 8073 6.61 33.33 8465 7.44
SPIRIT 88.55 1118 0.23 64.20 2144 0.57 46.67 7198 5.78 43.33 7817 6.57
ASAP 90.75 753 0.16 70.80 1649 0.43 46.67 5552 5.04 36.67 5434 5.10
Table 2: Experimental results of different methods on mathematical reasoning benchmarks with DeepSeek-R1-Distill-Qwen-7B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds. The best results are highlighted in bold, and the second-best are underlined.
4.1 Experimental Setup
Models and Datasets.
All experiments are conducted on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B Guo et al. (2025), with DeepSeek-R1-Distill-Qwen-7B as the default backbone across all settings. For the code reasoning domains, we use the Python subset of the CodeForces-CoTs Hugging Face (2025) dataset. For the math reasoning domain, we adopt the OpenR1-Math Hugging Face (2025) dataset and randomly sample 10K instances to match the size of the code subset, ensuring a balanced comparison across domains. The datasets consist of high-quality Chain-of-Thought (CoT) samples generated by DeepSeek-R1, making it particularly suitable for training competitive reasoning tasks. Detailed implementation settings (hyperparameters, hardware, etc.) are provided in Appendix C.
Benchmarks.
We evaluate our method on a suite of widely used benchmarks that cover both code generation and mathematical reasoning tasks. For code generation, we adopt HumanEval+ Chen et al. (2021); Liu et al. (2023), LiveCodeBench v1_v3, LiveCodeBench v4_v5 Jain et al. (2024), and LeetCodeDataset Xia et al. (2025b). For mathematical reasoning, we evaluate on GSM8K Cobbe et al. (2021), MATH500 Hendrycks et al. (2021), AIME24, and AIME25.
Baselines.
We compare our method against a comprehensive set of baselines. Zero-shot refers to the original model without any task-specific fine-tuning. Original denotes the model fine-tuned on the uncompressed CoTs from the training data. Among compression approaches, Selective Context Li et al. (2023) prunes redundant lexical units based on self-information; LLMLingua-2 Pan et al. (2024) distills GPT-4’s token importance signals into a lightweight Transformer encoder trained as a token classifier; TokenSkip Xia et al. (2025a) learns to skip less informative tokens to achieve controllable compression; and SPIRIT Cui et al. (2025) identifies critical reasoning steps by measuring perplexity shifts. Except for the zero-shot setting, all methods involve fine-tuning on CoTs processed according to their respective compression strategies.
Metrics.
We evaluate both accuracy and inference efficiency of each approach across three metrics: Pass@1 (Acc), which measures the percentage of problems correctly solved on the first attempt; Tokens (Tok), which denotes the average number of tokens generated by the LRMs; and Latency (Lat), which measures the average time (in seconds) required for the model generation.
4.2 Main Results
Tables 1 and 2 present the results of various methods on all benchmarks. The results show that the model fine-tuned on CoTs pruned by ASAP consistently achieves the best trade-off between accuracy and efficiency. It achieves the best accuracy while generating the fewest tokens, leading to the lowest generation latency.
We notice a clear distinction between token-level and step-level pruning strategies. Token-level baselines such as Selective Context, LLMLingua-2, and TokenSkip exhibit a significant performance degradation compared to the original CoTs. This is because the token removal disrupts the syntactic structure and semantic coherence of the original reasoning steps. Consequently, the fine-tuning data becomes fragmented and grammatically unnatural, making it difficult for the model to learn the intended logical flow of the CoT. Step-level methods, such as SPIRIT, perform significantly better than token-level pruning methods, due to the preservation of sentence-level integrity. While SPIRIT improves efficiency over the Original with comparable accuracy, our method achieves higher efficiency and accuracy at the same time. This improvement is particularly pronounced on the challenging LiveCodeBench v4_v5 benchmark: ASAP reduces the average number of generated tokens by 23.5% (from 7892 to 6035) and lowers generation latency by 43.5% (from 4.62s to 2.61s), while also achieving a 7.8% improvement in accuracy (Pass@1 increases from 33.58% to 36.19%).
4.3 Ablation and Analysis
Effect of Different Components.
To validate the contribution of each component, we conduct an ablation study on three model variants. 1) w/o Anchor-guided Pruning: which skips Stage 1 and applies only surprisal-based pruning to the original CoT. 2) w/o Surprisal-based Refining: which omits the surprisal-based refinement stage; and 3) w/o Both Pruning: equivalent to the original baseline, where the model is fine-tuned on the full, uncompressed CoT. Table 3 presents the results on LiveCodeBench v4_v5, which is representative of the consistent trends observed across benchmarks. Additional results are included in Appendix D. The results show that both pruning stages are essential and mutually complementary for optimal accuracy and efficiency. First, removing the anchor-guided pruning leads to a drop in both accuracy and efficiency. While the accuracy decrease is modest, the generation latency increases by a substantial 76.2% (from 2.61s to 4.60s), underscoring the importance of stage 1. Second, removing the surprisal-based refining results in a significant degradation across all metrics. The accuracy drops by 12.4% (Pass@1 decreases from 36.19% to 31.72%) relative to the ASAP, and efficiency improvements are largely lost. This highlights that our surprisal-based pruning mechanism is essential to select the most critical steps.
Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
ASAP 36.19 6035 2.61
w/o Anchor-guided Pruning 35.07 7735 4.60
w/o Surprisal-based Refining 31.72 8061 4.83
w/o Both Pruning 30.97 8289 4.83
Table 3: Ablation study of different pruning strategies on LiveCodeBench v4_v5. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Generalization to Different Architectures.
To validate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B. Following the same experimental protocol, we compare ASAP against three strong baselines: Zero-shot, Original, and SPIRIT. We observe consistent trends across all benchmarks. For brevity, we present representative results on two key benchmarks: LiveCodeBench v4_v5 and AIME24 in Table 4, while reporting the full results in the Appendix E. The results in the Llama3.1 series are highly consistent with our findings in the Qwen2.5 series, confirming the generalization of the ASAP. As shown in Table 4, ASAP achieves the highest accuracy on both benchmarks, and the efficiency improvements are even more pronounced. On LiveCodeBench, for instance, ASAP not only surpasses the accuracy of the Original baseline (32.84% vs. 31.34%) but also generates 49.1% fewer tokens and reduces latency by over 3x (from 8.60s to 2.69s). This suggests that the ASAP is particularly effective in identifying and distilling the core reasoning patterns, validating its robustness and broad applicability for improving reasoning efficiency across different model families.
Methods LCB AIME
Zero-shot 25.00 8508 8.90 33.33 8445 10.42
Original 31.34 8202 8.60 36.67 8550 10.04
SPIRIT 30.22 7913 8.45 36.67 8788 10.04
ASAP 32.84 4175 2.69 36.67 5314 6.97
Impact of Token Budget.
To evaluate the scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens to the model). We compare ASAP against the three strong baselines-SPIRIT, Original, and Zero-shot—across all benchmarks, and observe consistent trends. For clarity, we present results on LiveCodeBench v4_v5 under six budget settings ranging from 2K to 12K tokens. Results for other benchmarks and additional statistics are provided in Appendix F. As shown in Figure 4, ASAP consistently outperforms all baselines across all budget settings. In particular, ASAP exhibits smooth performance scaling with respect to the token budget. We note that ASAP achieves superior performance-efficiency trade-offs. For example, ASAP with just an 8K token budget achieves higher accuracy than SPIRIT and Original at a much larger 12K budget. These results further validate the practical utility of ASAP in real-world scenarios.
Training Efficiency.
To quantify the training efficiency gains, we present results of the CodeForces-CoTs dataset in Table 5 and results on other datasets are provided in Appendix G. The results highlight the training efficiency advantage of the ASAP. By generating the most compact yet logically rich CoTs, our approach significantly reduces training overhead. Compared to the uncompressed baseline (Original), our method reduces the number of training tokens by 75.6% and shortens training time by 60.7%. These savings substantially exceed those achieved by all other baselines. ASAP enables a more resource-efficient training process, making it a practical and cost-effective solution for real-world deployment.
Methods Tokens Time
Original 13023 80.11
Selective Context 6722 (-48.4%) 63.41 (-20.9%)
LLMLingua-2 6919 (-46.9%) 65.25 (-18.6%)
TokenSkip 9813 (-24.6%) 77.27 (-3.6%)
SPIRIT 6082 (-53.3%) 57.45 (-28.3%)
ASAP 3178 (-75.6%) 31.48 (-60.7%)
Table 5: Training efficiency comparison on CodeForces-CoTs dataset. We report the average number of tokens per sample and training time measured in seconds per step. Percentages indicate the reduction relative to the Original baseline.
5 Related Work
Chain-of-Thought and Advanced Reasoning.
Chain-of-Thought (CoT) prompting Wei et al. (2022) has evolved from heuristic prompting strategies Yao et al. (2023); Lei et al. (2023); Ling et al. (2023) to the training of specialized Large Reasoning Models (LRMs) like OpenAI’s o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025). These models leverage reinforcement learning to scale test-time compute, generating lengthy reasoning traces to solve complex tasks Kimi et al. (2025); Yang et al. (2025a); Yu et al. (2025); Wang et al. (2025); Zhang et al. (2025a, d). Unlike prior works that enhance performance by scaling up CoT length, we focus on pruning redundancy to improve efficiency without compromising reasoning performance.
Context Compression for LLMs.
To mitigate the computational cost of long contexts, various compression techniques have been proposed Zhang et al. (2025e). Approaches like Selective Context Li et al. (2023), LLMLingua series Jiang et al. (2023); Pan et al. (2024), and LongCodeZip Shi et al. (2025) employ information-theoretic metrics or small external models to filter redundant tokens. However, these methods typically treat input as unstructured text. Applying them directly to CoT often disrupts the syntactic and logical coherence required for valid reasoning, a limitation that our pruning aims to overcome.
Efficient Reasoning via Fine-Tuning.
Recent research has explored various efficiency mechanisms Qu et al. (2025), ranging from compressing thoughts into continuous latent representations Hao et al. (2024); Cheng and Van Durme (2024); Shen et al. (2025) to compressing CoTs Kang et al. (2025); Xia et al. (2025a); Cui et al. (2025). Approaches like TokenSkip Xia et al. (2025a) and SPIRIT Cui et al. (2025) reduce length by filtering tokens or steps based on heuristics or perplexity shifts. However, these metrics often struggle to differentiate between syntactic fluency and logical necessity. ASAP differs by combining anchor-guided structural pruning with first-token surprisal, offering a more robust proxy for cognitive pivots.
6 Conclusion
In this paper, we address the inefficiency of Large Reasoning Models stemming from the structural and logical redundancies in Chain-of-Thought reasoning. Grounded in an information-theoretic perspective, our large-scale empirical analysis reveals a fundamental property of reasoning traces: Information Concentration, where the logical uncertainty is highly concentrated at the onset of reasoning steps. Guided by this insight, we propose ASAP. This coarse-to-fine framework first aligns the reasoning structure with a logical anchor and then refines it using a novel First-Token Surprisal metric. Extensive experiments across multiple benchmarks demonstrate that ASAP outperforms existing baselines, establishing a new state-of-the-art Pareto frontier between accuracy and efficiency. Our work highlights the potential of using information-theoretic signals for efficient reasoning. Future work will explore applying ASAP to online inference acceleration.
Limitations
While ASAP demonstrates significant improvements in reasoning efficiency, we acknowledge several limitations. First, our method relies on the availability of a capable LLM to generate high-quality “Direct Thoughts” in Stage 1. If the anchor contains logical errors or hallucinations, it may misguide the subsequent pruning, although our pattern-matching constraint mitigates this risk. Second, our experiments primarily focus on code generation and mathematical reasoning. While we believe the principle of information concentration applies broadly, the effectiveness of ASAP on creative writing or commonsense reasoning tasks remains to be verified.
Appendix A Algorithms